OpenStudy (anonymous):

find the x-intercepts by factoring y=x^2-4x-32

sam (.sam.):

x^2-4x-32 Factor (x+4)(x-8)=0 set each brackets =0 (x+4)=0 , (x-8)=0

sam (.sam.):

you there?

OpenStudy (anonymous):

yeah,